Susan wants to mix 10lbs of virginia peanuts that cost $3.50 a pound with spanish peanuts that cost $3.00 a pound to obtain a mixture that cost $3.40 a pound. How many pounds of spanish peanuts should she use?
Answer by kapilsinghi(68)

virginia peanuts =10 lbs
cost of virginia peanuts = $3.5
total cost of virginia Peanuts = $35
let the spanish peanuts be x lbs
cost of spanish peanuts = $3
total cost of spanish Peanuts = $3x
total quantity = 10 + x
average cost of totol quantity = $3.4
total cost =(10+x)(3.4)=cost of virginia Peanuts + cost of spanish Peanuts
(10+x)(3.4)= 35 + 3x
34 + 3.4x = 35 + 3x
3.4x - 3x = 35 - 34
.4x = 1
x = 1/0.4
x = 10/4
x = 2.5
2.5 lbs
kapilsinghi123@gmail.com